# RadioDiff-FS: Physics-Informed Manifold Alignment in Few-Shot Diffusion Models for High-Fidelity Radio Map Construction

### 19 Mar 2026 v1

@@ IV-B Feature Shift Geometry
$U_{\mathrm{MU}}(\bm{x})=U_{\mathrm{MP}}(\bm{x})+\Delta U(\bm{x}),$ (33)
$\Delta U(\bm{x})\geq 0,\ \forall\bm{x}\in\Omega.$ (34)
To connect this physical structure with learning, introduce a fixed feature encoder
@@
which maps a RM field to a $d$-dimensional representation. The encoder $\Phi$ is kept fixed during adaptation and is used only to measure domain shift and structural consistency. We assume that the representation varies stably with the RM field.
@@
$\|\Phi(U_{\mathrm{MU}})-\Phi(U_{\mathrm{MP}})\|_{2}$ $=\|\Phi(U_{\mathrm{MP}}+\Delta U)-\Phi(U_{\mathrm{MP}})\|_{2},$
@@
Therefore, if $\Delta U$ is spatially localized or energy-limited as established in Section IV-A1, then the induced feature shift remains controlled, which supports stable transfer from MP-RM to MU-RM.
Let $U^{\mathrm{MP}}$ and $U^{\mathrm{MU}}$ be random fields representing the MP-RM and MU-RM domains under matched transmitter and environment semantics. Define
@@
The average shift induced by multipath is defined as
@@
Using (34) and (38), (40) admits the equivalent representation
@@
which shows that $\bm{w}$ is the mean feature displacement caused by the multipath residual.
The structural implication of Section IV-A1 is that the MP-RM skeleton is preserved while $\Delta U$ injects localized increments. In a representation that captures global geometry and coarse power layout, this motivates a translation dominant approximation in the feature space. Consider paired samples $\{(U_{i}^{\mathrm{MP}},U_{i}^{\mathrm{MU}})\}_{i=1}^{N}$ that share the same underlying configuration, and define
@@
$\mathbb{E}[\bm{\delta}_{i}]=\bm{0},$ (44)
$\|\bm{\delta}_{i}\|_{2}\leq\eta.$ (45)
The term $\bm{\delta}_{i}$ captures sample-specific multipath variability that is not explained by the common direction $\bm{w}$.
A first-order justification of (45) follows from local linearization. Assume $\Phi$ is Fréchet differentiable in a neighborhood of $U_{i}^{\mathrm{MP}}$ and write
@@
where $\mathcal{J}_{\Phi}(U)$ is the Fréchet derivative as a bounded linear operator and $\bm{r}_{i}$ is the remainder. Taking expectations and using (41) yields
$\bm{w}=\mathbb{E}\!\left[\mathcal{J}_{\Phi}(U^{\mathrm{MP}})[\Delta U]\right]+\mathbb{E}[\bm{r}],$ (47)
and (45) follows by identifying $\bm{\delta}_{i}$ with deviations of $\mathcal{J}_{\Phi}(U_{i}^{\mathrm{MP}})[\Delta U_{i}]+\bm{r}_{i}$ from its mean.
Under (45), pairwise relations are stable up to the fluctuation level $\eta$. For any $i,j$,
@@
Equation (49) formalizes that the global relational geometry induced by MP-RM features is largely preserved in the MU-RM domain, while the dominant inter-domain difference is captured by the common direction $\bm{w}$.
Given paired samples (42), define empirical means
$\widehat{\bm{\mu}}_{\mathrm{MP}}=\frac{1}{N}\sum_{i=1}^{N}\bm{z}_{i}^{\mathrm{MP}},$ (50)
$\widehat{\bm{\mu}}_{\mathrm{MU}}=\frac{1}{N}\sum_{i=1}^{N}\bm{z}_{i}^{\mathrm{MU}},$ (51)
and the empirical shift direction
@@
Equation (54) shows that the shift direction can be reliably estimated from a small number of paired samples when the translation dominant approximation holds with bounded fluctuations.
@@ IV-C Directional Consistency Principle
Let $\mathcal{G}_{\theta}$ denote a conditional generator that maps a MP-RM input $U^{\mathrm{MP}}$ and condition $c$ to a predicted MU-RM output
@@
Let $\Phi$ be the fixed feature encoder defined in (35), and define the feature displacement induced by the generator
@@
Let $\bm{w}=\bm{\mu}_{\mathrm{MU}}-\bm{\mu}_{\mathrm{MP}}$ be the mean shift direction defined in (40). The central requirement is that the generator update should align with the dominant cross-domain direction, while allowing sample-dependent magnitudes that reflect spatially varying multipath increments in RM fields.
@@
where $\alpha$ determines the amount of multipath injection and $\bm{e}$ represents orthogonal distortion. Directional consistency requires suppressing the orthogonal component $\bm{e}$ in (57).
To enforce this principle, decompose $\Delta\bm{z}$ into its projection onto $\bm{w}$ and its orthogonal residual. Define
@@
where $\beta\geq 0$ controls the penalty on negative alignment.
The first term in (63) enforces that the feature displacement lies predominantly in the one-dimensional subspace spanned by $\bm{w}$ and thus suppresses orthogonal distortions. The second term enforces nonnegative alignment, which is consistent with the nonnegative multipath increment property in the power domain established in Section IV-A1.
@@
Consequently, a generator can satisfy distance constraints while rotating the feature configuration, which changes the semantic axes along which MP-RM structure is represented and can distort the inherited skeleton. In contrast, minimizing (63) suppresses the orthogonal component (60) for every sample, which fixes the cross-domain displacement direction and eliminates the rotational ambiguity.
The directional constraint also yields a stability property consistent with Section IV-A2. Under the translation dominant model, a directionally consistent generator satisfies
$\Delta\bm{z}_{i}=\alpha_{i}\bm{w}+\bm{e}_{i},$ (65)
$\bm{e}_{i}^{\top}\bm{w}=0,$ (66)
$\|\bm{e}_{i}\|_{2}\leq\varepsilon.$ (67)
Then, for any $i,j$,
@@
where $\bm{z}_{i}^{\mathrm{gen}}=\Phi(\widehat{U}_{i}^{\mathrm{MU}})$. If $\bm{e}_{i}$ is controlled by the first term in (63), the deformation of the relational geometry is confined to a one-dimensional modulation along $\bm{w}$, which preserves the dominant MP-RM skeleton encoded by directions orthogonal to $\bm{w}$ and allocates the cross-domain change to the physically meaningful axis of multipath injection.
@@
where $\lambda_{\mathrm{dir}}>0$ controls the strength of the proposed principle. The regularizer in (69) steers adaptation along the mean shift direction of MP-RM to MU-RM while suppressing orthogonal distortions, thereby aligning cross-domain statistics without sacrificing the geometric integrity of the RM skeleton.
@@ IV-D Few-Shot Fine-Tuning
The objective is to adapt a pretrained generative model from an MP-RM domain to a MU-RM domain using only a small number of labeled MU-RM samples. The MP-RM domain provides abundant supervision that captures stable large-scale spatial priors and global propagation skeletons, while the MU-RM domain exhibits additional spatially localized power increments induced by multipath interactions and material heterogeneity. The adaptation should preserve the pretrained global structure and inject MU-RM specific details in a controlled and physically consistent manner.
Let $c$ denote the conditioning input that encodes scene geometry and transmitter attributes. The fine-tuning set is composed of paired samples
@@
where $N$ is small and each pair shares the same configuration specified by $c_{i}$.
As is shown in Fig. 2, we employ a variance-preserving diffusion model to parameterize the conditional generator. Let $x_{0}$ denote a clean MU-RM sample and let $q(x_{t}\mid x_{0})$ be the forward noising process indexed by $t\in[0,1]$. The perturbed sample is written as
@@
where $\alpha(t)$ is a monotone decreasing signal factor and $\sigma(t)$ is a monotone increasing noise factor. The denoiser predicts the injected noise $\varepsilon$ from $(x_{t},t,c)$. To allow different propagation components to be modeled with separate heads when needed, we write the network output as $\varepsilon_{\theta,r}(x_{t},t,c)$ for $r\in\mathcal{R}$ and define $\varepsilon_{r}$ as the corresponding ground-truth noise used in (71). The base objective is the denoising score matching loss
@@
To guide few-shot adaptation toward physically consistent updates, we introduce a directional constraint in a frozen semantic feature space. Let $\phi:\mathbb{R}^{H\times W}\rightarrow\mathbb{R}^{d}$ be a fixed encoder that maps a RM to a $d$ dimensional embedding. Define the mean embeddings of the MP-RM and MU-RM domains as
$\bm{\mu}_{\mathrm{MP}}=\mathbb{E}\!\left[\phi(x_{0}^{\mathrm{MP}})\right],\qquad\bm{\mu}_{\mathrm{MU}}=\mathbb{E}\!\left[\phi(x_{0}^{\mathrm{MU}})\right],$ (73)
and define the mean shift vector and its unit direction as
@@
For each triplet $(x_{0}^{\mathrm{MP}},x_{0}^{\mathrm{MU}},c)\in\mathcal{D}_{\mathrm{ft}}$, we define the target-aligned shift magnitude in feature space as
@@
Let $\widehat{x}_{0,\theta}(x_{t},t,c)$ denote the predicted clean sample produced by the denoiser at time $t$. Define the feature displacement relative to the MP-RM reference as
@@
Directional regularization is most effective when the denoised prediction $\widehat{x}_{0,\theta}$ is informative. We therefore weight $\mathcal{L}_{\mathrm{dir}}$ by a schedule that increases with the signal-to-noise ratio. Define the signal-to-noise ratio as
@@
and define a normalized increasing weight
@@
This objective preserves the pretrained MP-RM structural priors through $\mathcal{L}_{\mathrm{base}}$ while enforcing propagation-consistent feature shifts through $\mathcal{L}_{\mathrm{dir}}$, enabling data-efficient transfer to MU-RM with limited labeled samples.
TABLE I: Quantitative Comparison on IRT4. Results in bold red and underlined blue highlight the highest and second highest, respectively. “Rate vs. RadioDiff (%)” denotes the percentage improvement of the proposed method (RadioDiff-FS) compared to the baseline RadioDiff.
@@
PSNR $\uparrow$ 36.37 31.77 31.40 30.44 29.74 29.64 $\uparrow$ 19.48%
TABLE II: Quantitative Comparison on IRT4 With Car. Results in bold red and underlined blue highlight the highest and second highest, respectively. “Rate vs. RadioDiff (%)” denotes the percentage improvement of the proposed method (RadioDiff-FS) compared to the baseline RadioDiff.
@@ V-A Datasets
Figure 3: The comparisons of constructed SRM on different methods.
We evaluate on RadioMapSeer [32], a large-scale urban benchmark for RM prediction. Each scene is represented by a $256\times 256$ building occupancy grid with $1$ m spatial resolution and a unified building height of $25$ m. We consider the device-to-device (D2D) setting with Tx and Rx heights of $1.5$ m, a carrier frequency of $5.9$ GHz, and a Tx power of $23$ dBm [32] RadioMapSeer provides subsets with different propagation complexities. We use the multipath-rich subset with up to four interactions, denoted as IRT4 in the dataset release, as the target MU-RM domain. A key characteristic of IRT4 is its limited supervision density: each scene contains only two transmitter placements, hence only two MU-RM samples per scene. This results in $1402$ MU-RM samples over $701$ scenes, which naturally forms a few-shot adaptation setting when fine-tuning a pretrained generator. We evaluate both static RMs (SRM) that account for building-induced attenuation and dynamic RMs (DRM) that additionally include vehicle-induced blockage [32]. We adopt a scene-disjoint split with $500$ scenes for training and validation and $200$ unseen scenes for testing, ensuring generalization to new urban layouts. Unless otherwise stated, few-shot fine-tuning uses only a small subset of the IRT4 training samples, while the remaining samples are used for validation and ablation.
V-B Metrics and Baseline
@@
Figure 4: The comparisons of constructed DRM on different methods.
We compare RadioDiff-FS with representative baselines covering supervised regression, adversarial generation, and diffusion-based generation. To isolate the effect of the proposed few-shot adaptation, all methods are trained from scratch on the same few-shot supervision from the IRT4 subset, using $500$ training scenes with $2$ transmitter placements per scene, and are evaluated on the identical scene-disjoint test split.
@@
    RadioUNet [10]: A sampling-free convolutional regressor based on a U-Net backbone that predicts RMs directly from the environment and transmitter inputs.
@@
    RME-GAN [35]: An adversarial generator adapted to the sampling-free protocol by using only environmental features and transmitter metadata, enabling direct RM synthesis without sparse measurements.
@@
RadioDiff is particularly relevant because it belongs to the same backbone family as our approach and directly maps the environment and transmitter inputs to the RM. It therefore serves as a controlled ablation baseline that removes the proposed main-path to multipath guidance mechanism while keeping the generative capacity unchanged.

### 26 Mar 2026 v3

@@ IV-B Feature Shift Geometry
$U_{\mathrm{MU}}(\bm{x})=U_{\mathrm{MP}}(\bm{x})+\Delta U(\bm{x}),\quad\Delta U(\bm{x})\geq 0,\ \forall\,\bm{x}\in\Omega.$ (33)
To connect this physical structure with learning, we introduce a fixed feature encoder
@@
which maps a RM field to a $d$-dimensional representation. The encoder $\Phi$ is kept fixed during adaptation and serves only to measure domain shift and structural consistency. We assume that the representation varies stably with the RM field, as stated below.
@@
$\|\Phi(U_{\mathrm{MU}})-\Phi(U_{\mathrm{MP}})\|_{2}$ $=\|\Phi(U_{\mathrm{MP}}+\Delta U)-\Phi(U_{\mathrm{MP}})\|_{2}$
@@
If $\Delta U$ is spatially localized or energy-limited as established in Section IV-A, the induced feature shift remains controlled. This result supports stable transfer from MP-RM to MU-RM in the learned representation.
Let $U^{\mathrm{MP}}$ and $U^{\mathrm{MU}}$ be random fields representing the MP-RM and MU-RM domains under matched transmitter and environment semantics. Define the corresponding feature vectors as
@@
The average feature shift induced by multipath is defined as
@@
Using (33) and (37), the vector $\bm{w}$ in (39) admits the equivalent representation
@@
which shows that $\bm{w}$ is the mean feature displacement caused by the multipath residual. The structural implication of Section IV-A is that the MP-RM skeleton is preserved while $\Delta U$ injects localized increments. In a representation that captures global geometry and coarse power layout, this observation motivates a translation-dominant approximation in the feature space.
Consider paired samples $\{(U_{i}^{\mathrm{MP}},U_{i}^{\mathrm{MU}})\}_{i=1}^{N}$ sharing the same underlying configuration, and define
@@
$\mathbb{E}[\bm{\delta}_{i}]=\bm{0},\quad\|\bm{\delta}_{i}\|_{2}\leq\eta,$ (43)
where $\bm{\delta}_{i}$ captures sample-specific multipath variability not explained by the common direction $\bm{w}$.
A first-order justification of (42) follows from local linearization. Assume that $\Phi$ is Fréchet differentiable in a neighborhood of $U_{i}^{\mathrm{MP}}$, and write
@@
where $\mathcal{J}_{\Phi}(U)$ denotes the Fréchet derivative as a bounded linear operator and $\bm{r}_{i}$ is the higher-order remainder. Taking expectations and using (40) yields
$\bm{w}=\mathbb{E}\!\left[\mathcal{J}_{\Phi}(U^{\mathrm{MP}})[\Delta U]\right]+\mathbb{E}[\bm{r}].$ (45)
The translation model (42) then follows by identifying $\bm{\delta}_{i}$ with the deviation of $\mathcal{J}_{\Phi}(U_{i}^{\mathrm{MP}})[\Delta U_{i}]+\bm{r}_{i}$ from its mean.
Under (42), pairwise relations are stable up to the fluctuation level $\eta$. Then, for any $i$ and $j$, we have
@@
Equation (47) formalizes that the relational geometry of MP-RM features is largely preserved in the MU-RM domain. The dominant inter-domain difference is captured by the common shift direction $\bm{w}$.
Given the paired features in (41), define empirical means as
$\widehat{\bm{\mu}}_{\mathrm{MP}}=\frac{1}{N}\sum_{i=1}^{N}\bm{z}_{i}^{\mathrm{MP}},\qquad\widehat{\bm{\mu}}_{\mathrm{MU}}=\frac{1}{N}\sum_{i=1}^{N}\bm{z}_{i}^{\mathrm{MU}},$ (48)
and define the empirical shift direction as
@@
Equation (51) shows that the shift direction can be reliably estimated from a small number of paired samples. This guarantee holds when the translation-dominant approximation is valid with bounded fluctuations.
@@ IV-C Directional Consistency Principle
Let $\mathcal{G}_{\theta}$ denote a conditional generator that maps an MP-RM input $U^{\mathrm{MP}}$ and condition $c$ to a predicted MU-RM output
@@
Let $\Phi$ be the fixed feature encoder defined in (34). The feature displacement induced by the generator is defined as
@@
Let $\bm{w}=\bm{\mu}_{\mathrm{MU}}-\bm{\mu}_{\mathrm{MP}}$ be the mean shift direction defined in (39). The central requirement is that the generator update should align with the dominant cross-domain direction. At the same time, the alignment should allow sample-dependent magnitudes that reflect spatially varying multipath increments.
@@
where $\alpha$ determines the amount of multipath injection and $\bm{e}$ represents orthogonal distortion. Directional consistency requires suppressing the orthogonal component $\bm{e}$.
To enforce this principle, we decompose $\Delta\bm{z}$ into its projection onto $\bm{w}$ and its orthogonal residual. Define
@@
where $\beta\geq 0$ controls the penalty on negative alignment. The first term in (60) enforces that the feature displacement lies in the one-dimensional subspace spanned by $\bm{w}$, thereby suppressing orthogonal distortions. The second term enforces nonnegative alignment, which is consistent with the nonnegative multipath increment property established in Section IV-A.
@@
A generator can therefore satisfy distance constraints while rotating the feature configuration. Such rotation changes the semantic axes along which the MP-RM structure is represented and can distort the inherited skeleton. In contrast, minimizing (60) suppresses the orthogonal component for every sample. This fixes the cross-domain displacement direction and eliminates the rotational ambiguity.
The directional constraint also yields a stability property consistent with Section IV-B. Under the translation-dominant model, a directionally consistent generator satisfies
$\Delta\bm{z}_{i}=\alpha_{i}\,\bm{w}+\bm{e}_{i},\quad\bm{e}_{i}^{\top}\bm{w}=0,\quad\|\bm{e}_{i}\|_{2}\leq\varepsilon.$ (62)
Then, for any $i$ and $j$, we have
@@
where $\bm{z}_{i}^{\mathrm{gen}}=\Phi(\widehat{U}_{i}^{\mathrm{MU}})$. When $\bm{e}_{i}$ is controlled by the first term in (60), the deformation of the relational geometry is confined to a one-dimensional modulation along $\bm{w}$. This preserves the dominant MP-RM skeleton encoded by directions orthogonal to $\bm{w}$ and allocates the cross-domain change to the physically meaningful axis of multipath injection.
@@
where $\lambda_{\mathrm{dir}}>0$ controls the regularization strength. This regularizer steers adaptation along the mean shift direction from MP-RM to MU-RM while suppressing orthogonal distortions. It thereby aligns cross-domain statistics without sacrificing the geometric integrity of the RM skeleton.
@@ IV-D Few-Shot Fine-Tuning
The objective is to adapt a pretrained generative model from the MP-RM domain to the MU-RM domain using only a small number of labeled MU-RM samples. The MP-RM domain provides abundant supervision that captures stable large-scale spatial priors and global propagation skeletons. In contrast, the MU-RM domain exhibits additional spatially localized power increments induced by multipath interactions. The adaptation should preserve the pretrained global structure and inject MU-RM-specific details in a controlled and physically consistent manner. Let $c$ denote the conditioning input that encodes scene geometry and transmitter attributes. The fine-tuning set is composed of paired samples
@@
where NN is small and each pair shares the same configuration specified by $c_{i}$.
As shown in Fig. 2, we employ a variance-preserving diffusion model to parameterize the conditional generator. Let $x_{0}$ denote a clean MU-RM sample and let $q(x_{t}\mid x_{0})$ be the forward noising process indexed by $t\in[0,1]$. The perturbed sample is written as
@@
where $\alpha(t)$ is a monotonically decreasing signal factor and $\sigma(t)$ is a monotonically increasing noise factor. The denoiser predicts the injected noise $\varepsilon$ from the tuple $(x_{t},t,c)$. To allow different propagation components to be modeled with separate heads, we write the network output as $\varepsilon_{\theta,r}(x_{t},t,c)$ for $r\in\mathcal{R}$ and define $\varepsilon_{r}$ as the corresponding ground-truth noise in (66). The base objective is the denoising score matching loss
@@
To guide few-shot adaptation toward physically consistent updates, we introduce a directional constraint in a frozen semantic feature space. Let $\phi:\mathbb{R}^{H\times W}\rightarrow\mathbb{R}^{d}$ be a fixed encoder that maps a RM to a $d$-dimensional embedding. The mean embeddings of the MP-RM and MU-RM domains are defined as
$\bm{\mu}_{\mathrm{MP}}=\mathbb{E}\!\left[\phi(x_{0}^{\mathrm{MP}})\right],\qquad\bm{\mu}_{\mathrm{MU}}=\mathbb{E}\!\left[\phi(x_{0}^{\mathrm{MU}})\right].$ (68)
The mean shift vector and its unit direction are given by
@@
For each triplet $(x_{0}^{\mathrm{MP}},x_{0}^{\mathrm{MU}},c)\in\mathcal{D}_{\mathrm{ft}}$, the target-aligned shift magnitude in feature space is defined as
@@
Let $\widehat{x}_{0,\theta}(x_{t},t,c)$ denote the predicted clean sample produced by the denoiser at diffusion time $t$. The feature displacement relative to the MP-RM reference is
@@
Directional regularization is most effective when the denoised prediction $\widehat{x}_{0,\theta}$ is informative. We therefore weight $\mathcal{L}_{\mathrm{dir}}$ by a schedule that increases with the signal-to-noise ratio (SNR). The SNR at diffusion time $t$ is defined as
@@
and the normalized increasing weight is given by
@@
This objective preserves the pretrained MP-RM structural priors through $\mathcal{L}_{\mathrm{base}}$ while enforcing propagation-consistent feature shifts through $\mathcal{L}_{\mathrm{dir}}$. It enables data-efficient transfer from the MP-RM domain to the MU-RM domain with limited labeled samples.
TABLE I: Quantitative Comparison on IRT4. Results in bold red and underlined blue highlight the best and second best, respectively. The last column denotes the percentage improvement of RadioDiff-FS over the best baseline without fine-tuning.
@@
PSNR $\uparrow$ 36.37 0.9285 31.40 30.44 29.74 29.64 $\uparrow$ 19.48%
TABLE II: Quantitative Comparison on IRT4 With Car. Results in bold red and underlined blue highlight the best and second best, respectively. The last column denotes the percentage improvement of RadioDiff-FS over the best baseline without fine-tuning.
@@ V-A Datasets
Figure 3: Visual comparison of constructed SRM across different methods.
We evaluate on RadioMapSeer [33], a large-scale urban benchmark for RM prediction. Each scene is represented by a $256\times 256$ building occupancy grid with 11 m spatial resolution and a unified building height of $25$ m. We consider the device-to-device (D2D) setting with transmitter and receiver heights of $1.5$ m, a carrier frequency of $5.9$ GHz, and a transmitter power of $23$ dBm [33]. RadioMapSeer provides subsets with different propagation complexities. We use the multipath-rich subset with up to four interactions, denoted as IRT4 in the dataset release, as the target MU-RM domain. A key characteristic of IRT4 is its limited supervision density: each scene contains only two transmitter placements and thus only two MU-RM samples per scene. This results in $1402$ MU-RM samples over $701$ scenes, which naturally forms a few-shot adaptation setting when fine-tuning a pretrained generator. We evaluate both static RMs (SRM) that account for building-induced attenuation and dynamic RMs (DRM) that additionally include vehicle-induced blockage [33]. We adopt a scene-disjoint split with $500$ scenes for training and validation and $200$ unseen scenes for testing, ensuring generalization to new urban layouts. Unless otherwise stated, few-shot fine-tuning uses only a small subset of the IRT4 training samples, while the remaining samples are used for validation and ablation.
V-B Metrics and Baselines
@@
Figure 4: Visual comparison of constructed DRM across different methods.
We compare RadioDiff-FS with representative baselines covering supervised regression, adversarial generation, and diffusion-based generation. To isolate the effect of the proposed few-shot adaptation, all methods are trained from scratch on the same few-shot supervision from the IRT4 subset. Specifically, all methods use $500$ training scenes with 22 transmitter placements per scene, and are evaluated on the identical scene-disjoint test split. We additionally report RadioDiff-FS in a one-shot setting, where only a single MU-RM sample per scene is used for fine-tuning, to examine performance under extreme data scarcity. The baselines are summarized as follows.
@@
    RadioUNet [10]: A sampling-free convolutional regressor based on a U-Net backbone that predicts RMs directly from environment and transmitter inputs.
@@
    RME-GAN [35]: An adversarial generator adapted to the sampling-free protocol by using only environmental features and transmitter metadata for direct RM synthesis.
@@
RadioDiff is particularly relevant because it belongs to the same backbone family as our approach. It directly maps the environment and transmitter inputs to the RM without the proposed main-path to multipath guidance mechanism. It therefore serves as a controlled ablation baseline that isolates the contribution of the directional consistency constraint while keeping the generative capacity unchanged.